OpenStudy (anonymous):

solve for x: 25^(x+1)*125^(x)=25^(3x-2). Please show me the steps

OpenStudy (anonymous):

5^(2x+2+3x)=5^(6x-4) hence, 5x+2=6x-4 so, x=6

OpenStudy (anonymous):

Thanks